# GLProtein: Global-and-Local Structure Aware Protein Representation Learning

### v1

@@ Abstract
Proteins are central to biological systems, participating as building blocks across all forms of life. Despite advancements in understanding protein functions through protein sequence analysis, there remains potential for further exploration in integrating protein structural information. We argue that the structural information of proteins is not only limited to their 3D information but also encompasses information from amino acid molecules (local information) to protein-protein structure similarity (global information). To address this, we propose GLProtein, the first framework in protein pre-training that incorporates both global structural similarity and local amino acid details to enhance prediction accuracy and functional insights. GLProtein innovatively combines protein-masked modelling with triplet structure similarity scoring, protein 3D distance encoding and substructure-based amino acid molecule encoding. Experimental results demonstrate that GLProtein outperforms previous methods in several bioinformatics tasks, including predicting protein-protein interaction, contact prediction, and so on.
@@ 1 Introduction
Proteins are fundamental to virtually every biological process, serving as the building blocks for cells and organs and acting as catalysts, messengers, and structural elements in all life forms. Understanding the structure and function of proteins is crucial for advances in health guo2017roles ; williams2019plasma ; li2022dietary ; ding2019selective , agriculture rischer2020cellular ; peydayesh2021protein ; davis2024food , and environmental science wheeler2021environmental ; sobhi2023advanced ; zhao2024holistic ; weng2021late , making protein research a cornerstone of biotechnology and medicinal science. Recognizing the critical role of proteins in various scientific fields, many efforts have been made to design computational methods to further understand these crucial molecules sliwoski2014computational ; zhao2020exploring ; li2024tomg . Particularly, protein representation learning, as one significant part, involves capturing the complex features and relationships within proteins in a condensed form that can be utilized for various computational tasks and analyses. It is crucial for enhancing the understanding of protein structures and functions, improving predictive modelling in bioinformatics, facilitating the drug discovery process, and advancing our knowledge of biological systems through interpretable and efficient representations of proteins somnath2021multi ; liu2023generative ; gao2024drugclip .
@@
In recent years, the success of language models zhao2024recommender ; hu2023improving ; li2024empowering ; li2024molreflect in natural language processing (NLP) has paved the way for innovative approaches in bioinformatics areas, such as protein modeling xiao2021modeling ; chowdhury2022single , protein generation madani2020progen ; ferruz2022protgpt2 , and protein-protein interaction prediction wang2019high ; ofer2021language . To be specific, by treating protein sequences as linguistic strings, these models have demonstrated remarkable effectiveness in predicting protein function based on sequence data alone. Technically, as shown in Figure 1, protein sequences (e.g., the amino acid sequence ‘MLTAHV…’) are treated as sentences in natural language and amino acids (e.g., ‘M’, ‘L’, and ‘T’) resemble words. Thus, Leveraging the powerful BERT architecture originally developed for natural language, ProtBert elnaggar2021prottrans adeptly adapts the BERT devlin2018bert masked language modelling framework to the field of bioinformatics. This analogy allows ProtBert to employ the technique of predicting randomly masked elements in sequences, thereby learning to identify complex patterns and dependencies among amino acids. Similar to ProtBert, ESM rives2021biological ; verkuil2022language ; hie2022high extends this paradigm by employing a more refined Transformer-based architecture, focusing on capturing the evolutionary relationships and functional dynamics within protein sequences. In other words, most existing protein modelling methods aim to perform protein representation learning by encoding the protein’s sequence information for various downstream applications, such as amino acid contact prediction singh2022spot , protein homology detection kaminski2023plm , protein stability prediction chu2024protein , protein-protein interaction identification wang2019high ; ofer2021language , etc.
@@
Despite the aforementioned successes, most existing protein language modelling methods suffer from intrinsic limitations. Specifically, most of their focuses have primarily been on the amino acid sequence, often neglecting the crucial aspects of protein structure. Proteins possess the ability to fold into diverse 3D shapes, interacting with various proteins and small molecules in biologically significant ways jumper2021highly ; mirdita2022colabfold ; tsaban2022harnessing . Since protein’s structure determines function greslehner2018molecular , utilizing protein 3D structure information effectively is crucial for protein language modelling, in which many studies have demonstrated the potential of pre-training on experimentally determined protein structures hermosilla2022contrastive ; su2023saprot ; wang2022lm ; zhang2022protein . Nevertheless, these methods focus only on the structure within proteins and ignore the global similarities between proteins. We emphasize that the information on protein structure is not only limited to its structure (i.e., conformation) in 3D space but also includes information ranging from local amino acid molecules to the global structural similarity between proteins, as shown in Figure 1. Local information involves the detailed properties and orientations of individual amino acids, which can affect protein stability and biochemical activity renaud2021deeprank . These specifics are vital as they demonstrate how modifications or mutations at the amino acid level can alter the overall structure and functionality of the protein jumper2021highly . Furthermore, protein structure similarities provide information on evolutionary relationships and functional classes, which are crucial for understanding how structurally similar proteins of different species can perform similar or complementary functions within biological systems hamamsy2023protein . For example, as shown in Figure 2, the bacterial ice-binding protein FfIBP and the diatom adhesion protein CaTrailin_4 exhibit no detectable sequence similarity despite their functional similarities zackova2023diatom ; al2023protein . Their predicted structures exhibit a remarkable similarity (TM-Score = 0.6), with both proteins adopting a beta-helical fold comprised of two units linked by an alpha helix. This structural topology is characteristic of ice-binding proteins. Such comparisons are key to predicting the functions of newly discovered proteins based on known structures, thereby enhancing our grasp of complex biological processes and interactions lipman1985rapid ; hamamsy2022tm ; hamamsy2023protein . However, most existing approaches have ineffectively incorporated amino acid molecule information and protein structural similarities into protein representation learning.
@@
(c) Structure Alignment
Figure 2: An example of protein structure similarity. Given the predictive structures of a protein pair: (a) the bacterial ice-binding protein FfIBP and (b) the diatom adhesion protein CaTrailin_4 zackova2023diatom ; al2023protein , (c) is FfIBP (blue) and CaTrailin_4 (green) structure alignment.
@@ 2 Related Work
In this section, we briefly review related work about protein language modelling as well as protein structure modelling.
@@
Protein Langauge Modeling. As an approach to protein representation learning, protein language modelling is a burgeoning field at the intersection of computational biology and natural language processing (NLP). Inspired by the success of language models in NLP, researchers have adapted these techniques to analyze and predict the properties of protein sequences fan2025computational . Recent advancements have been dominated by the application of transformer-based models, which utilize self-attention mechanisms to capture relationships between amino acids in a sequence. ProtTrans elnaggar2021prottrans and ESM beal2015esm ; verkuil2022language ; hie2022high , trained on large-scale protein databases, have shown remarkable ability in tasks such as protein classification and interaction prediction. Moreover, OntoProtein zhangontoprotein and KeAP zhou2023protein incorporated external biological knowledge to enrich protein representations and enhance performance on various downstream tasks. However, most of these protein language models do not explicitly consider the spatial structure of proteins and structural similarities between proteins like our proposed approach.
@@
Figure 3: Overview of our proposed model, which jointly optimizes global protein similarities and masked protein model with local structure information.
@@
Protein Structure Modelling. The structure of a protein determines its functions. Thus, protein structure modelling has been treated as a reliable way to improve protein representation learning huang2024protein ; alquraishi2021machine ; torrisi2020deep ; cheng2008machine . Some methods use Graph Neural Networks (GNNs) to handle the complex, non-linear relationships inherent in protein structure jha2022prediction ; reau2023deeprank ; xu2024deeprank . Moreover, RGN2 chowdhury2022single utilized a protein language model to learn structural information from unaligned protein sequences. GearNet zhang2022protein focused on geometric pertaining and learned protein features by utilizing spatial relationships between amino acids. SaProt su2023saprot introduced the concept of a "structure-aware vocabulary" to integrate residue tokens with structure tokens. Similar to the knowledge hancing method, PST chen2024endowing enhances protein language models by integrating structural information through graph transformers to incorporate structural data. Unlike these models, we propose global structure learning and local structure learning methods, which could not only integrate protein structure information and amino acid information but also learn the structure similarity between different proteins by using TM-Score hamamsy2023protein .
3 The Proposed Method GLProtein
@@
In this section, we aim to introduce our proposed framework (GLProtein) as a novel solution to learn global and local protein structure information for protein representation learning.
3.1 An Overview of the Proposed Framework
Most existing protein language modelling methods often have limitations, primarily concentrating on linear sequence information while overlooking the essential aspects of protein structure. To address such limitations, we develop GLProtein, a novel protein pre-training framework that incorporates both global and local protein structure information into protein representation learning. More specifically, in order to integrate structure information effectively, a global protein triplet contrastive learning, protein 3D distance encoding, and substructure-based molecular encoding fusion are introduced to empower the protein language model for protein representation learning. The framework of GLProtein, shown in Figure 3, consists of three components: protein language modelling, global structure information modelling, and local structure information modelling. The first component is protein language modelling, which is the backbone of our proposed framework. It employs random masking to the amino acid sequence and treats local structure information as associated structure information, including 3D distance encoding and substructure-based molecular encoding. The second component is global structure information modelling, as input data in this component is still protein sequence. To obtain a protein triplet, we apply positive and negative protein sampling. After that, we use self-supervised contrastive learning to calculate the protein triplet loss and learn the protein similarity with global structure information by taking advantage of the huge amount of unlabeled data. The third component is local structure information modelling, which integrates local protein structure information in the protein decoder, including 3D distance encoding and substructure-based molecular encoding. Next, we will detail each model component.
@@ 3.2 Protein Language Modelling
Drawing inspiration from the success of masked language modelling in natural language processing, masked protein modelling conceals part of a protein sequence and tasks a model with predicting the hidden residues devlin2018bert ; elnaggar2021prottrans . As shown in the center part of Figure  3, protein language modelling forms the backbone of our proposed framework, which aims to learn protein representation. We adopt a masking strategy similar to KeAP zhou2023protein , where 20% of the amino acids in a sequence are randomly selected for masking. Each masked amino acid has an 80% probability of being masked for prediction, a 10% chance of being replaced with a random amino acid, and a 10% chance of remaining unchanged. We then integrate protein 3D distance encoding and substructure-based molecular encoding into a protein decoder, in which we will detail in the local structure information modelling component. Suppose that the number of masked amino acids is $M$ and $x_{i}$ denotes the $i$-th amino acid. $x_{\sim i}$ denotes the sequence of amino acids excluding the masked amino acid at position $i$. We leverage a cross-entropy loss $\mathcal{L}_{MLM}$ to estimate masked amino acids. Formally, the masked protein modelling objective can be defined as:
@@
where $\theta_{E}$ and $\theta_{D}$ denote the parameters of the protein sequence encoder and decoder, respectively. We initialize with a pre-trained BERT-like encoder: ProtBert elnaggar2021prottrans . Next, we will introduce the global structure information modelling part.
@@ 3.3 Global Structure Information Modelling
In protein science, proteins follow the sequence-structure-function paradigm koehler2023sequence , where the amino acid sequence dictates the protein’s three-dimensional structure, ultimately determining its function. Therefore, effectively utilizing information about protein 3D structures is essential for improving protein language modelling. However, protein structures encompass more than mere 3D spatial configurations; they also include global structural information that reflects similarities among proteins. Understanding these structural similarities is crucial, as they can provide insights into evolutionary relationships and functional roles, allowing for more accurate predictions of protein behaviour and interactions hamamsy2023protein . To address this, we introduce the concept of global structure information, which contains the structure similarities between proteins, by leveraging of the huge amount of self-supervised signals in protein sequences, as shown in the top of the Figure 3. To be specific, given for each input protein sequence, positive and negative protein sampling is designed to get the triplet $(P,P_{pos},P_{neg})$ for capturing protein structure similarity features. Then, the protein triplets are encoded to protein representation for the calculation of the contrastive learning loss. This optimizes the protein sequence encoder by bringing the representation of the input protein $P$ and its positive samples $P_{pos}$ closer together while pushing the representation of $P$ and its negative samples $P_{neg}$ further apart in the representation space.
3.3.1 Positive and Negative Protein Sampling
TM-score (Template Modeling Score) zhang2004scoring ; xu2010significant is a widely used metric in structural biology for assessing the structural similarity between two protein structures. One of its key advantages is the ability to effectively identify similar three-dimensional structures, even among proteins with low sequence similarity. Additionally, the TM-score is more sensitive to global fold similarity than to local structural variations, allowing it to capture essential information about the overall architecture of proteins. Consequently, we utilize the TM-score to measure structural similarity between proteins, focusing on their overall global structure rather than mere sequence identity. Mathematically, the TM-score can be expressed as:
@@ 3.3.1 Positive and Negative Protein Sampling
We employ a two-pronged approach that utilizes the TM-Vec model hamamsy2023protein to construct a robust set of positive and negative samples for our protein structure similarity analysis. For positive sample selection, we utilize the TM-Vec model to identify the top-K protein sequences that exhibit the highest TM-score values in relation to the template proteins. This approach ensures that our positive samples share significant structural similarity with the template.
In contrast, our negative sampling strategy employs a stochastic selection process followed by structural dissimilarity confirmation. Initially, we randomly select $n$ proteins from our dataset. Subsequently, we employ the TM-Vec model to compute the TM-score between each selected protein and the template protein. Proteins with a TM-score < 0.2 are classified as negative samples, as this threshold indicates a high degree of structural dissimilarity xu2010significant . This method ensures that our negative samples are structurally distinct from the template, providing a clear contrast to the positive samples.
3.3.2 Protein Triplet Modelling
To refine the embedding space by ensuring that proteins with similar structures or functions are clustered closer together than those that are dissimilar, we need to specify an objective function to optimize. Since the task we focus on in this part is contrastive learning, the protein triplet loss is designed. This loss function operates by comparing three entities: anchor protein $P$, positive protein $P_{pos}$ and negative protein $P_{neg}$. Thus, given protein representation triplet $(P,P_{pos},P_{neg})$, the protein triplet loss $\mathcal{L}_{PTL}$ can be defined as:
$\begin{split}\mathcal{L}_{PTL}(P,P_{pos},P_{neg})=max(||\mathbf{E}_{p}-\mathbf{E}^{pos}_{p}||_{2}-||\mathbf{E}_{p}-\mathbf{E}^{neg}_{p}||_{2}+\epsilon,0),\end{split}$ (3)

### v2

@@ Abstract
Proteins are central to biological systems, participating as building blocks across all forms of life. Despite advancements in understanding protein functions through protein sequence analysis, there remains potential for further exploration in integrating protein structural information. We argue that the structural information of proteins is not only limited to their 3D information but also encompasses information from amino acid molecules (local information) to protein-protein structure similarity (global information). To address this, we propose GLProtein, the first framework in protein pre-training that incorporates both global structural similarity and local amino acid details to enhance prediction accuracy and functional insights. GLProtein innovatively combines protein-masked modelling with triplet structure similarity scoring, protein 3D distance encoding and substructure-based amino acid molecule encoding. Experimental results demonstrate that GLProtein outperforms previous methods in several bioinformatics tasks, including predicting protein-protein interactions, contact prediction, and so on.
GLProtein: Global-and-Local Structure Aware Protein Representation Learning
Yunqing Liu     Wenqi Fanthanks: Corresponding author† ‡     Xiao-Yong Wei† §     Qing Li Department of Computing (COMP), The Hong Kong Polytechnic University Department of Management and Marketing (MM), The Hong Kong Polytechnic University §Department of Computer Science, Sichuan University yunqing617.liu@connect.polyu.hk wenqifan03@gmail.com, {cs007.wei, csqli}@comp.polyu.edu.hk
@@ 1 Introduction
Proteins are fundamental to virtually every biological process, serving as the building blocks for cells and organs and acting as catalysts, messengers, and structural elements in all life forms. Understanding the structure and function of proteins is crucial for advances in health, agriculture, and environmental science, making protein research a cornerstone of biotechnology and medicinal science Ding et al. (2019); Davis et al. (2024); Zhao et al. (2024). Recognizing the critical role of proteins in various scientific fields, many efforts have been made to design computational methods to further understand these crucial molecules Weng et al. (2021); Zhao et al. (2020). Particularly, protein representation learning, as one significant part, involves capturing the complex features and relationships within proteins in a condensed form that can be utilized for various computational tasks and analyses. It is crucial for enhancing the understanding of protein structures and functions, improving predictive modelling in bioinformatics, facilitating the drug discovery process, and advancing our knowledge of biological systems through interpretable and efficient representations of proteins Somnath et al. (2021); Liu et al. (2023); Gao et al. (2024).
@@
In recent years, the success of language models in natural language processing (NLP) has paved the way for innovative approaches in bioinformatics areas, such as protein modeling Xiao et al. (2021); Chowdhury et al. (2022), protein generation Madani et al. (2020); Ferruz et al. (2022), and protein-protein interaction prediction Wang et al. (2019); Ofer et al. (2021). To be specific, by treating protein sequences as linguistic strings, these models have demonstrated remarkable effectiveness in predicting protein function based on sequence data alone. Technically, as shown in Figure 1, protein sequences (e.g., the amino acid sequence ‘MLTAHV…’) are treated as sentences in natural language and amino acids (e.g., ‘M’, ‘L’, and ‘T’) resemble words. Thus, Leveraging the powerful BERT architecture originally developed for natural language, ProtBert Elnaggar et al. (2021) adeptly adapts the BERT Devlin et al. (2018) masked language modelling framework to the field of bioinformatics. This analogy allows ProtBert to employ the technique of predicting randomly masked elements in sequences, thereby learning to identify complex patterns and dependencies among amino acids. Similar to ProtBert, ESM Rives et al. (2021); Verkuil et al. (2022); Hie et al. (2022) extends this paradigm by employing a more refined Transformer-based architecture, focusing on capturing the evolutionary relationships and functional dynamics within protein sequences. In other words, most existing protein modelling methods aim to perform protein representation learning by encoding the protein’s sequence information for various downstream applications, such as amino acid contact prediction Singh et al. (2022), protein homology detection Kaminski et al. (2023), protein stability prediction Chu et al. (2024), protein-protein interaction identification Wang et al. (2019); Ofer et al. (2021), etc.
@@
Despite the aforementioned successes, most existing protein language modelling methods suffer from intrinsic limitations. Specifically, most of their focuses have primarily been on the amino acid sequence, often neglecting the crucial aspects of protein structure. Proteins possess the ability to fold into diverse 3D shapes, interacting with various proteins and small molecules in biologically significant ways Jumper et al. (2021); Mirdita et al. (2022); Tsaban et al. (2022). Since protein’s structure determines function Greslehner (2018), utilizing protein 3D structure information effectively is crucial for protein language modelling, in which many studies have demonstrated the potential of pre-training on experimentally determined protein structures Hermosilla and Ropinski (2022); Su et al. (2023); Wang et al. (2022); Zhang et al. (2022). Nevertheless, these methods focus only on the structure within proteins and ignore the global similarities between proteins. We emphasize that the information on protein structure is not only limited to its structure (i.e., conformation) in 3D space but also includes information ranging from local amino acid molecules to the global structural similarity between proteins, as shown in Figure 1. Local information involves the detailed properties and orientations of individual amino acids, which can affect protein stability and biochemical activity Renaud et al. (2021). These specifics are vital as they demonstrate how modifications or mutations at the amino acid level can alter the overall structure and functionality of the protein Jumper et al. (2021). Furthermore, protein structure similarities provide information on evolutionary relationships and functional classes, which are crucial for understanding how structurally similar proteins of different species can perform similar or complementary functions within biological systems Hamamsy et al. (2023). For example, as shown in Figure 2, the bacterial ice-binding protein FfIBP and the diatom adhesion protein CaTrailin_4 exhibit no detectable sequence similarity despite their functional similarities Zackova Suchanova et al. (2023); Al-Fatlawi et al. (2023). Their predicted structures exhibit a remarkable similarity (TM-Score = 0.6), with both proteins adopting a beta-helical fold comprised of two units linked by an alpha helix. This structural topology is characteristic of ice-binding proteins. Such comparisons are key to predicting the functions of newly discovered proteins based on known structures, thereby enhancing our grasp of complex biological processes and interactions Lipman and Pearson (1985); Hamamsy et al. (2022, 2023). However, most existing approaches have ineffectively incorporated amino acid molecule information and protein structural similarities into protein representation learning.
@@
(c) Alignment
Figure 2: An example of protein structure similarity. Given the predictive structures of a protein pair: (a) the bacterial ice-binding protein FfIBP and (b) the diatom adhesion protein CaTrailin_4 Zackova Suchanova et al. (2023); Al-Fatlawi et al. (2023), (c) is FfIBP (blue) and CaTrailin_4 (green) structure alignment.
@@ 2 Related Work
Protein Langauge Modelling. As an approach to protein representation learning, protein language modelling is a burgeoning field at the intersection of computational biology and natural language processing (NLP) Hu et al. (2023); Li et al. (2024b, c, a). Inspired by the success of language models in NLP, researchers have adapted these techniques to analyse and predict the properties of protein sequences Fan et al. (2025). Recent advancements have been dominated by the application of transformer-based models, which utilise self-attention mechanisms to capture relationships between amino acids in a sequence. ProtTrans Elnaggar et al. (2021) and ESM Beal (2015); Verkuil et al. (2022); Hie et al. (2022), trained on large-scale protein databases, have shown remarkable ability in tasks such as protein classification and interaction prediction. Moreover, OntoProtein Zhang et al. and KeAP Zhou et al. (2023a) incorporated external biological knowledge to enrich protein representations and enhance performance on various downstream tasks. However, most of these protein language models do not explicitly consider the spatial structure of proteins and structural similarities between proteins, like our proposed approach.
@@
Figure 3: Overview of our proposed model, which jointly optimises global protein similarities and masked protein model with local structure information.
@@
Protein Structure Modelling. The structure of a protein determines its functions. Thus, protein structure modelling has been treated as a reliable way to improve protein representation learning Huang et al. (2024); AlQuraishi (2021); Torrisi et al. (2020); Cheng et al. (2008). Some methods use Graph Neural Networks (GNNs) to handle the complex, non-linear relationships inherent in protein structure Liao et al. (2023); Jha et al. (2022); Réau et al. (2023); Xu and Bonvin (2024); Zhou et al. (2023b, 2024); Li et al. (2025). Moreover, RGN2 Chowdhury et al. (2022) utilized a protein language model to learn structural information from unaligned protein sequences. GearNet Zhang et al. (2022) focused on geometric pertaining and learned protein features by utilizing spatial relationships between amino acids. SaProt Su et al. (2023) introduced the concept of a "structure-aware vocabulary" to integrate residue tokens with structure tokens. Similar to the knowledge hancing method, PST Chen et al. (2024) enhances protein language models by integrating structural information through graph transformers to incorporate structural data. Unlike these models, we propose global structure learning and local structure learning methods, which could not only integrate protein structure information and amino acid information but also learn the structure similarity between different proteins by using TM-Score Hamamsy et al. (2023).
3 Methodology
@@
In this section, we aim to introduce our proposed framework (GLProtein) as a novel solution to learn global and local protein structure information for protein representation learning. We develop GLProtein that incorporates both global and local protein structure information into protein representation learning. The framework of GLProtein, shown in Figure 3, consists of three components: protein language modelling (Section 3.1), global structure information modelling (Section 3.2), and local structure information modelling (Section 3.3).
@@ 3.1 Protein Language Modelling
As shown in the center part of Figure  3, protein language modelling forms the backbone of our proposed framework, which aims to learn protein representation. We adopt a masking strategy that each masked amino acid has an 80% probability of being masked for prediction, a 10% chance of being replaced with a random amino acid, and a 10% chance of remaining unchanged. We then integrate protein 3D distance encoding and substructure-based molecular encoding into a protein decoder, in which we will detail in the local structure information modelling component. Suppose that the number of masked amino acids is $M$ and $x_{i}$ denotes the $i$-th amino acid. $x_{\sim i}$ denotes the sequence of amino acids excluding the masked amino acid at position $i$. We leverage a cross-entropy loss $\mathcal{L}_{MLM}$ to estimate masked amino acids. Formally, the masked protein modelling objective can be defined as:
@@
where $\theta_{E}$ and $\theta_{D}$ denote the parameters of the protein sequence encoder and decoder, respectively. We initialise with a pre-trained BERT-like encoder: ProtBert Elnaggar et al. (2021).
@@ 3.2 Global Structure Information Modelling
Protein structures encompass more than mere 3D spatial configurations; they also include global structural information that reflects similarities among proteins. To address this, we introduce the concept of global structure information, which contains the structure similarities between proteins, by leveraging the huge amount of self-supervised signals in protein sequences, as shown at the top of Figure 3. To be specific, given each input protein sequence, positive and negative protein sampling is designed to get the triplet $(P,P_{pos},P_{neg})$ for capturing protein structure similarity features. Then, the protein triplets are encoded to protein representation for the calculation of the contrastive learning loss. This optimises the protein sequence encoder by bringing the representation of the input protein $P$ and its positive samples $P_{pos}$ closer together while pushing the representation of $P$ and its negative samples $P_{neg}$ further apart in the representation space.
Positive and Negative Protein Sampling.
TM-score (Template Modeling Score) Zhang and Skolnick (2004); Xu and Zhang (2010) is a widely used metric in structural biology for assessing the structural similarity between two protein structures. We utilize the TM-score to measure structural similarity between proteins, focusing on their overall global structure rather than mere sequence identity. Mathematically, the TM-score can be expressed as:
$\text{TM-score}=max[\frac{1}{L_{N}}\sum^{L_{r}}_{i=1}\frac{1}{1+(\frac{d_{i}}{d_{0}})^{2}}],$ (2)
We employ a two-pronged approach that utilizes the TM-Vec model Hamamsy et al. (2023) to construct a robust set of positive and negative samples for our protein structure similarity analysis. For positive sample selection, we utilize the TM-Vec model to identify the top-K protein sequences that exhibit the highest TM-score values in relation to the template proteins.
In contrast, our negative sampling strategy employs a stochastic selection process followed by structural dissimilarity confirmation. Initially, we randomly select $n$ proteins from our dataset. Subsequently, we employ the TM-Vec model to compute the TM-score between each selected protein and the template protein. Proteins with a TM-score < 0.2 are classified as negative samples, as this threshold indicates a high degree of structural dissimilarity Xu and Zhang (2010).
Protein Triplet Modelling.
Since the task we focus on in this part is contrastive learning, the protein triplet loss is designed. This loss function operates by comparing three entities: anchor protein $P$, positive protein $P_{pos}$ and negative protein $P_{neg}$. Thus, given protein representation triplet $(P,P_{pos},P_{neg})$, the protein triplet loss $\mathcal{L}_{PTL}$ can be defined as: